suppose the market for grass seed can be expressed as follows demand =100-2p and supply =3p at the equilibrium ,calculate the price elasticity of demand and supply
1
Expert's answer
2020-04-15T09:17:52-0400

With given supply and demand curve the equilibrium price and quantity are P=20. Q=60.

Point elasticity formula: 

"E=\frac{dQ}{dP}\times\frac{P}{Q}."

So, "Ed=(-2)\times\frac{20}{60}=-\frac{2}{3}" .

"Es = 3\times\frac{20}{60}=1."

The price elasticity of demand is -2/3, the price elasticity of supply is 1.